The sum of an AP is 20 the 1st term being 8 and the common difference is -2 find the number of terms in the sequence
Answer by mananth(16946)

the formula for sum of n terms =
Sn = n/2(2a+(n-1)d)
20=n/2(2*8+(n-1)*-2)
40 = n(16-2n+2)
40=16n-2n^2+2n
40=18n-2n^2
2n^2-18n+40=0
divide by 2
n^2-9n+20=0
n^2-5n-4n+20=0
n(n-5)-4(n-5)=0
(n-5)(n-4)=0
n= 5 OR 4